# Rethinking Whole-Body CT Image Interpretation: An Abnormality-Centric Approach

### v1

@@ 3.5 Summary
We prioritize axial CT images for annotation, while also incorporating coronal and sagittal images based on clinical recommendations. In total, we annotated 18,969 abnormalities on 9,990 axial scans, 2,738 coronal scans, and 1,803 sagittal scans. Representative examples are shown in Figure 1, and the distribution of annotated abnormalities is detailed in Appendix Section A.11. Annotation quality is verified by a senior radiologist with 12 years’ experience, with details in Appendix A.5.
4 OminiAbnorm-CT
This section presents the details of OminiAbnorm-CT, a novel system for grounded CT image interpretation . We start with the problem formulation in Section 4.1, then its architectural details in Section 4.2 and finally the training details in Section 4.3.
@@ 4.1 Problem Formulation
Given a CT image ($\mathcal{I}\in\mathbb{R}^{H\times W\times D}$), this paper aims to build a model $\Phi_{\theta}(\cdot)$, that enables generating the textual descriptions for abnormal regions, with intermediate groundings when necessary:
@@ 4.2 Architecture
Figure 3: OminiAbnorm-CT. We bridge a multi-modal language model and a segmentation module, to allow grounding evidence acquisition during the generation of abnormality description, and further enhance its comprehension for flexible usage with text instruction and visual prompts.
The overall architecture of OminiAbnorm-CT is illustrated in Figure 3, which comprises a standard multi-modal language model and a segmentation module. Our core innovation lies in the integration of them, which enables localizing the abnormality and generating a textual description (findings), supporting flexible interaction through both visual prompts and textual instructions.
Multi-modal language model functions as the core to comprehend user instruction, invoke the segmentation module properly, and document the abnormal findings based on the grounding evidence. Specifically, in one workflow to generate grounded reports, given a CT image ($\mathcal{I}$) and text instruction ($\mathcal{T}$) as user input, it will first reason for the specific abnormal findings of interest for the user (or simply any abnormalities). For instance, when referring to patient history of tuberculous empyema, it should focus on related abnormalities including pleural thickening, calcifications, etc.
@@
Given the segmentation results ($S$), we convert it to visual prompt ($\mathcal{V}$) in box format, i.e., directly overlay it on the original image to create a new input image ($\mathcal{I}^{\prime}$), and the multi-modal language model could seamlessly continue generating the abnormality description ($\mathcal{R}$) based on it:
@@
where $\oplus$ represents the superimposing operation. This leverages the inherent ability of multi-modal language model to perceive visual markers cai2024vip , guiding its attention to specific regions of interest.
Note that, the above design naturally supports another workflow when users have manually delineated specific abnormalities, or intend to refine the grounding results. In such case, $\mathcal{V}$ can also be user-input visual prompts on the image, and the multi-modal language model elaborates the user-highlighted abnormality following Equation 3.
@@
where $v$ denotes concatenation of multi-scale image embeddings along the channel dimension, which are down-sampled from each encoder layer to a unified resolution, and $u$ is the pixel-level dense feature from the last decoder layer. To bridge the gap between latent spaces, we further align the segmentation prompt ($h_{\text{seg}}$) with the image embeddings ($v$):
@@ 4.3 Training
Training object. We train the multi-modal language model and the segmentation module jointly with a text generation loss and a segmentation loss:
$\mathcal{L}_{\textbf{txt}}=\text{CE}(\hat{\mathcal{R}},\mathcal{R}),\text{\hskip 6.0pt}\mathcal{L}_{\textbf{seg}}=\text{BCE}(\hat{\mathcal{S}},\mathcal{S})+\text{DICE}(\hat{\mathcal{S}},\mathcal{S})$ (7)
Training data. Our training data primarily originates from OminiAbnorm-CT-14K, and is formulated into three instruction data formats as illustrated in Section 3.4. Additionally, we incorporate two types of data as supplements: (i) lesion segmentation data. This type of data can be formulated into grounding tasks and utilized to enhance the grounding ability. A detailed list of these datasets can be found in Appendix Table 6; (ii) medical VQA data. This type of data is involved to maintain our model’s generalization capabilities with more diverse question types. Specifically, we take CT images and corresponding question-answering data from PubMedVision chen2024huatuogpt .
Implementation details. In practice, we adopt QWen2.5-VL-7B yang2024qwen2 as the multi-modal language model, and a 6-layer U-Net as the segmentation module. The cross-attention module consists of 6 layers. As shown in Appendix Section A.8, we find that expanding the input context range (i.e., using more slices) does not enhance the report generation performance of the multi-modal language model. Thus, we only feed the central slice to the model. However, for the segmentation module, we include up to 8 adjacent slices as additional input. To save computational cost, we only optimize the inserted LoRA layers hu2022lora in the multi-modal language model while leaving the whole segmentation module trainable. More implementation details are in Appendix Section A.3.
@@ 5 Experiments
In this section, we first present experimental settings in Section 5.1, including task formulation, metrics, baselines, and test datasets. We then provide detailed quantitative experimental results in Section 5.2, and ablation study in Section 5.3. In addition, the qualitative experimental results are in the Appendix Section A.1 and more ablation studies are in the Appendix A.8 and A.9.
Model B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
@@
GPT-4o 15.80 6.15 1.59 39.04 84.58 19.31 18.62 14.41 6.51
QWen2.5-VL-7B 11.57 4.20 0.76 38.35 84.10 16.68 14.26 11.09 3.85
ViP-LLaVA 15.21 4.49 0.69 33.07 85.15 13.65 17.41 14.68 3.35
MedFlamingo ✓ 11.26 1.71 0.00 17.35 80.34 9.43 11.69 10.55 0.50
MedDr ✓ 9.54 2.46 0.41 32.24 78.80 9.13 12.23 10.34 3.74
LLaVA-Med ✓ 15.55 4.23 0.55 39.08 84.68 15.58 17.67 13.55 6.06
BiomedGPT ✓ 10.80 4.16 0.71 38.16 83.19 15.77 13.40 10.81 4.65
OminiAbnorm-CT ✓ 18.03 6.65 1.80 42.81 86.35 19.40 21.66 17.61 9.98
@@
GPT-4o 9.34 3.15 0.96 27.13 58.83 9.27 12.25 10.01 4.31
QWen2.5-VL-7B 13.62 3.54 0.66 37.26 84.19 13.49 14.97 11.47 3.37
ViP-LLaVA 13.85 3.81 0.53 35.21 84.60 12.52 17.07 13.96 3.80
MedFlamingo ✓ 10.98 1.66 0.00 17.73 79.98 9.01 11.60 10.34 0.53
MedDr ✓ 8.07 2.07 0.29 32.77 76.07 8.09 11.55 9.63 4.08
LLaVA-Med ✓ 15.65 4.18 0.68 40.99 84.38 14.18 18.15 13.52 6.54
BiomedGPT ✓ 12.60 4.79 0.70 38.47 82.89 15.45 14.78 11.52 4.45
OminiAbnorm-CT ✓ 18.38 7.19 2.39 42.88 86.00 19.00 21.30 16.87 9.44
@@
GPT-4o 10.29 3.40 0.84 32.90 72.76 10.07 13.86 11.63 4.04
QWen2.5-VL-7B 13.28 3.54 0.63 35.67 83.98 13.01 14.75 11.34 2.91
ViP-LLaVA 12.84 3.83 0.38 34.51 84.44 11.88 16.29 13.44 2.86
MedFlamingo ✓ 10.57 1.53 0.00 17.36 78.87 8.66 11.40 10.14 0.36
MedDr ✓ 6.98 1.77 0.28 30.77 75.35 7.12 10.32 8.54 2.42
LLaVA-Med ✓ 15.10 4.68 0.66 39.19 84.10 14.18 17.90 13.57 5.11
BiomedGPT ✓ 12.59 4.81 0.73 36.62 82.60 15.19 14.62 11.35 3.37
OminiAbnorm-CT ✓ 17.29 6.48 1.82 41.77 85.62 17.45 20.43 15.87 7.51
@@ 5.1 Experiment Settings
Based on OminiAbnorm-CT-14K, we build a rigorous and comprehensive benchmark to evaluate both grounding and description generation abilities for abnormalities on whole-body CT images.
@@
Metrics. For grounding results, we calculate Dice Similarity Coefficient (DSC); for generated descriptions, we calculate BLEU (B) bleu , RaTEScore (RTSc) zhao2024ratescore , BERTScore (BTSc) zhangbertscore , METEOR (MTR) banerjee2005meteor , ROUGE (R) rouge and RadGraph (RG) delbrouck2024radgraph . In the text-guided grounded report generation, we prompt models to output a specific response when the queried abnormalities are confirmed absent. Specificity (SP) is then calculated as the proportion of correctly identified absent abnormalities out of all truly absent abnormalities in the test set. Meanwhile, when models miss the queried abnormalities, both DSC and all description metrics are assigned as 0.
Baseline. For visual prompted report generation, we compare with multi-modal LLMs in both the medical domain (MedDr meddr , BiomedGPT zhang2024generalist , LLaVA-Med li2023llava , and Med-Flamingo moor2023med ) and the general domain (GPT-4o gpt4o , Qwen2.5-VL yang2024qwen2 , and ViP-LLaVA cai2024vip ). For (text-guided) grounded report generation, as no such solution exists in the literature, we prompt LLaVA-Med to describe abnormal findings based on segmentation results from auxiliary grounding models. Since the images in OminiAbnorm-CT-14K lack the original DICOM metadata, we only consider 2D segmentation models as baselines, including BiomedParse zhao2024foundation , LiSA lai2024lisa , and MedULS de2024uls23 . More details are in Appendix Section A.4.
Data. We independently split the train and test data for axial, coronal, and sagittal images. To guarantee diversity, we allocate at least 5 samples per category to the axial test set (2 for coronal and sagittal), or all samples if fewer exist, with the remaining data following a 3:1 train-test ratio. To avoid data leakage, images from the same patient are restricted to the same set.
5.2 Experiment Results
Model DSC B-1 B-2 B-3 RTSc BTSc MTR R-1 R-L RG
@@ 5.2 Experiment Results
LiSA + LLaVA-Med 20.96 12.94 6.65 0.73 31.38 82.07 19.20 14.62 12.44 2.79
BiomedParse + LLaVA-Med 15.78 13.61 6.98 0.84 31.50 82.36 19.84 15.28 13.02 2.74
MedULS + LLaVA-Med 15.24 12.70 6.43 0.74 31.05 84.27 18.74 14.67 12.42 2.31
OminiAbnorm-CT 36.04 18.50 11.75 7.90 33.24 86.10 21.83 21.65 18.75 3.78
@@
LiSA + LLaVA-Med 17.09 11.48 5.85 0.64 33.28 82.27 16.23 14.36 12.12 2.74
BiomedParse + LLaVA-Med 14.55 11.71 5.95 0.68 32.89 81.80 16.88 14.75 12.57 3.35
MedULS + LLaVA-Med 12.77 11.64 5.88 0.64 32.25 84.11 16.45 14.71 12.34 2.96
OminiAbnorm-CT 31.65 19.13 11.86 7.81 35.47 85.58 20.66 22.14 18.26 4.55
@@
LiSA + LLaVA-Med 14.72 11.10 5.60 0.58 31.51 81.12 16.18 14.07 11.96 2.18
BiomedParse + LLaVA-Med 14.45 10.89 5.46 0.52 32.10 79.37 15.68 14.06 11.95 2.13
MedULS + LLaVA-Med 12.95 10.95 5.32 0.50 30.94 83.53 15.39 14.12 11.83 1.87
OminiAbnorm-CT 34.38 17.10 10.63 7.10 33.13 85.25 19.05 20.92 17.67 3.70
Visual Prompted Report Generation. As depicted in Table 1, OminiAbnorm-CT consistently outperforms all baselines across all planes and metrics. Specifically, compared to the strongest baseline LLaVA-Med, OminiAbnorm-CT achieves average improvements of 2.5 in BLEU-1, 2.7 in RaTEScore, and 3.1 in RadGraph across all three planes. These results demonstrate OminiAbnorm-CT’s effectiveness in semi-automated workflows, where radiologists identify abnormalities and the model provides accurate and detailed interpretations.
Grounded Report Generation. As shown in Table 2, OminiAbnorm-CT demonstrates superior grounding capabilities, outperforming the best baselines by 15.08, 14.56, and 19.66 in DSC across axial, coronal, and sagittal planes respectively. For report generation, OminiAbnorm-CT achieves the best performance across all metrics, with notable improvements in both RadGraph and RaTEScore across all anatomical planes, indicating that the generated reports can better align with clinical standards. These results demonstrate OminiAbnorm-CT’s potential for fully automated abnormality localization and report generation in clinical settings.

### v2

@@ 4.4 Summary
We prioritize axial images for annotation, while also incorporate coronal and sagittal images based on clinical recommendations. In total, we annotated 18,969 abnormalities on 9,990 axial images, 2,738 coronal images, and 1,803 sagittal images. Representative examples and the distribution of annotated abnormalities are shown in Figure 1. Annotation quality and inter-annotator consistency is verified by a radiologist with 12-year experience, detailed in the Appendix.
5 OmniAbnorm-CT
This section presents the details of OmniAbnorm-CT, a novel system for grounded CT image interpretation . We start with the problem formulation in Section 5.1, then its architectural details in Section 5.2 and finally the training details in Section 5.3.
Figure 3: OmniAbnorm-CT. We bridge a VLM and a segmentation module, to allow grounding evidence acquisition during the generation of abnormality description, and further enhance its comprehension for flexible usage with text instruction and visual prompts.
@@ 5.1 Problem Formulation
Given a CT image ($\mathcal{I}\in\mathbb{R}^{H\times W\times D}$), we aims to build a model $\Phi_{\theta}(\cdot)$, that generates the textual descriptions for abnormalities, with intermediate groundings when necessary:
@@
Model Rubric BLEU-1 BLEU-2 BLEU-3 RaTESc BERTSc METEOR ROUGE-1 ROUGE-L RadG
Axial (n=2193)
GPT4o 33.41 15.90 6.22 1.63 39.04 84.57 19.61 18.78 14.55 6.52
QWen25VL7B(OG) 7.14 5.19 0.98 0.13 14.38 37.14 5.86 5.92 4.47 0.72
ViP-LLaVA 8.31 15.24 4.50 0.63 33.07 85.12 13.67 17.49 14.69 3.28
MedDr ✓ 20.82 9.43 2.35 0.33 32.27 78.62 9.01 12.19 10.29 3.67
LLaVA-Med ✓ 15.41 15.62 4.25 0.55 39.11 84.65 15.72 17.76 13.60 6.13
BiomedGPT ✓ 20.37 10.85 4.25 0.72 38.29 83.18 15.88 13.50 10.91 4.64
OmniAbnorm-CT ✓ 35.69 18.03 6.65 1.80 42.81 86.35 19.40 21.66 17.61 9.98
Coronal (n=750)
GPT4o 25.41 9.25 3.12 0.92 27.26 59.17 9.25 12.08 9.90 4.26
QWen25VL7B(OG) 26.25 13.43 3.43 0.65 37.34 84.19 13.34 14.73 11.27 3.34
ViP-LLaVA 10.27 13.69 3.67 0.51 35.08 84.58 12.54 16.91 13.88 3.68
MedDr ✓ 26.79 8.22 2.03 0.26 32.98 76.53 8.17 11.71 9.81 4.13
LLaVA-Med ✓ 22.34 15.60 4.11 0.65 40.96 84.39 14.15 17.99 13.45 6.46
BiomedGPT ✓ 20.25 12.58 4.79 0.74 38.69 82.90 15.37 14.69 11.46 4.39
OmniAbnorm-CT ✓ 31.30 18.38 7.19 2.39 42.88 86.00 19.00 21.30 16.87 9.44
Sagittal (n=591)
GPT4o 24.39 10.23 3.45 0.88 32.41 72.53 10.01 13.77 11.65 4.00
QWen25VL7B(OG) 20.09 13.22 3.47 0.59 35.51 83.93 12.82 14.54 11.23 2.95
ViP-LLaVA 11.56 12.87 3.81 0.38 34.24 84.39 11.86 16.19 13.44 2.84
MedDr ✓ 19.79 6.89 1.70 0.26 30.32 74.72 6.91 10.07 8.37 2.29
LLaVA-Med ✓ 15.68 15.01 4.66 0.69 38.97 84.06 14.13 17.64 13.46 5.09
BiomedGPT ✓ 17.59 12.37 4.73 0.70 36.24 82.54 14.93 14.35 11.22 3.12
OmniAbnorm-CT ✓ 29.00 17.29 6.48 1.82 41.77 85.62 17.45 20.43 15.87 7.51
@@ 5.2 Architecture
The overall architecture of OmniAbnorm-CT is illustrated in Figure 3, which comprises a VLM and a segmentation module. Our core innovation lies in the integration of them, which enables localizing the abnormality and generating a textual description (findings), supporting flexible interaction through both visual prompts and textual instructions.
Vision language model functions as the core to comprehend user instruction, invoke the segmentation module properly, and document the abnormal findings based on the grounding evidence. Specifically, in one workflow to generate grounded reports, given a CT image ($\mathcal{I}$) and text instruction ($\mathcal{T}$) as user input, it will first reason for the specific abnormal findings of interest for the user (or simply any abnormalities). For instance, when referring to patient history of tuberculous empyema, it should focus on related abnormalities including pleural thickening, calcifications, etc.
@@
Given the segmentation results ($S$), we convert it to visual prompt ($\mathcal{V}$) in box format, i.e., directly overlay it on the original image to create a new input image ($\mathcal{I}^{\prime}$), and the VLM could seamlessly continue generating the abnormality description ($\mathcal{R}$) based on it:
@@
where $\oplus$ represents superimposition. This leverages the inherent ability of VLM to perceive visual markers [cai2024vip], guiding its attention to specific regions of interest.
Note that, the above design naturally supports another workflow when users have manually delineated specific abnormalities, or intend to refine the grounding results. In such case, $\mathcal{V}$ can also be user-input visual prompts on the image, and the VLM elaborates the user-highlighted abnormality following Equation 3.
@@
where $v$ denotes concatenation of multi-scale image embeddings along the channel dimension, which are down-sampled from each encoder layer to a unified resolution, and $u$ is the pixel-level dense feature from the last decoder layer. To bridge the gap between latent spaces, we align the segmentation prompt ($h_{\text{seg}}$) with the image embeddings ($v$):
@@ 5.3 Training
Training object. We train the VLM and the segmentation module jointly with a text generation loss and a segmentation loss:
[…]
$\mathcal{L}_{\textbf{seg}}$ $=\text{BCE}(\hat{\mathcal{S}},\mathcal{S})+\text{DICE}(\hat{\mathcal{S}},\mathcal{S})$ (8)
Training data. Our training data primarily originates from OmniAbnorm-CT-14K, and is formulated into three instruction data formats as illustrated in Section 4.3. Additionally, we incorporate two types of data as supplements: (i) lesion segmentation data. This type of data can be formulated into grounding tasks and utilized to enhance the grounding ability. A detailed list of these datasets can be found in the Appendix; (ii) medical visual question-answering (VQA) data. This type of data is involved to maintain our model’s generalization capabilities with more diverse question types. Specifically, we take CT images and corresponding question-answering data from PubMedVision [chen2024huatuogpt].
Implementation details. In practice, we adopt QWen2.5-VL-7B [yang2024qwen2] as the VLM, and a 6-layer U-Net as the segmentation module. The cross-attention module consists of 6 layers. We feed 9 consecutive slices to the segmentation module, matching the annotation setup. Our collaborating radiologists confirm this window provides sufficient local 3D context for the vast majority of clinically significant findings. However, as detailed in the Appendix, expanding the input context (i.e., multiple slices) does not improve VLM report generation. We therefore feed only the center slice to the VLM. To save computational cost, we only optimize the inserted LoRA layers [hu2022lora] in the VLM while leaving the whole segmentation module trainable. More details, runtime and computational analysis are in the Appendix.
@@ 6 Experiments
Model DSC Rubric B-1 B-2 B-3 RaTESc BERTSc MTR R-1 R-L RadG
@@
MedULS + LLaVA-Med 14.92 3.00 10.60 5.31 0.50 26.84 72.92 15.90 12.46 10.50 2.26
LiSA + LLaVA-Med 20.62 4.87 12.85 6.53 0.70 31.09 81.68 19.11 14.46 12.32 2.75
BiomedParse + LLaVA-Med 15.69 3.45 13.35 6.86 0.80 31.61 83.04 19.62 14.98 12.77 2.84
OmniAbnorm-CT ✓ 36.04 13.03 18.50 11.75 7.90 33.24 86.10 21.83 21.65 18.75 3.78
@@
MedULS + LLaVA-Med 12.73 4.03 9.54 4.91 0.58 28.65 72.77 13.75 12.61 10.68 2.24
LiSA + LLaVA-Med 17.61 5.66 10.76 5.54 0.68 33.04 81.89 16.04 14.20 11.91 2.94
BiomedParse + LLaVA-Med 14.77 6.53 11.94 6.07 0.67 32.35 81.29 16.73 15.18 12.64 3.06
OmniAbnorm-CT ✓ 31.65 13.29 19.13 11.86 7.81 35.47 85.58 20.66 22.14 18.26 4.55
@@
MedULS + LLaVA-Med 12.87 2.16 9.71 4.97 0.70 26.06 69.51 13.18 12.26 10.50 1.64
LiSA + LLaVA-Med 14.82 4.36 12.10 6.16 0.57 31.40 82.77 16.95 14.84 12.76 2.00
BiomedParse + LLaVA-Med 13.72 4.51 12.25 6.21 0.57 31.90 81.70 16.93 14.98 12.82 2.31
OmniAbnorm-CT ✓ 34.38 10.11 17.10 10.63 7.10 33.13 85.25 19.05 20.92 17.67 3.70
In this section, we first illustrate the experimental settings in Section 6.1, including test data, task formulation, and baselines; Then we introduce the evaluation metrics for different tasks in Section 6.2, including existing metrics, and propose AbnormRubric for report generation evaluation; Finally, we provide detailed internal validation results in Section 6.3, and external validation results in Section 6.4. Due to limited space, qualitative experiment results, and ablation studies on the segmentation module, context range, and visual prompts are presented in the Appendix.
@@ 6.1 Experiment Settings
Internal validation data. We split OmniAbnorm-CT-14K into the train and test set for training and internal validation, following three principles: (i) Axial, coronal, and sagittal images are split independently; (ii) To guarantee a comprehensive benchmark, we allocate at least 5 samples per category to the axial test set (2 for coronal and sagittal), or all samples if fewer exist, with the remaining data following a 3:1 train-test ratio. (iii) To avoid data leakage, images from the same patient are restricted to the same set.
External validation data. In addition, we apply the annotation pipeline described in Section 4.2 to 65 chest CT scans sampled from CT-RATE [hamamci2024developing], a large-scale non-contrast chest CT dataset collected in Turkey clinical centers, which serves as an external validation set to assess generalizability. Annotators are instructed to label all visible abnormalities in the full volume and, for each abnormality, to select the most prominent slice (i.e., the slice with the largest segmentation area) for evaluation. This yields a total of 271 abnormality annotations.
@@
Baseline. For visual prompted report generation, we compare with multi-modal LLMs in both the medical domain (MedDr [meddr], BiomedGPT [zhang2024generalist], and LLaVA-Med [li2023llava]) and the general domain (GPT-4o [gpt4o], Qwen2.5-VL [yang2024qwen2], and ViP-LLaVA [cai2024vip]). For (text-guided) grounded report generation, as no such solution exists in the literature, we prompt LLaVA-Med to describe abnormal findings based on segmentation results from auxiliary grounding models. Since the images in OmniAbnorm-CT-14K lack the original DICOM metadata, we only consider 2D segmentation models as baselines, including BiomedParse [zhao2024foundation], LiSA [lai2024lisa], and MedULS [de2024uls23]. More details are in the Appendix.
Model DSC Rubric B-1 B-2 B-3 RaTESc BERTSc MTR R-1 R-L RadG
@@
MedULS + LLaVA-Med 14.24 3.58 9.27 2.32 0.27 26.58 78.54 11.06 11.03 8.86 1.78
LiSA + LLaVA-Med 17.34 0.25 1.78 0.41 0.25 4.40 22.11 2.03 2.34 2.08 0.24
BiomedParse + LLaVA-Med 16.42 0.56 1.91 0.44 0.19 4.73 23.34 2.15 2.54 2.22 0.27
OmniAbnorm-CT ✓ 32.40 14.82 11.94 3.33 0.74 34.80 83.85 12.94 14.65 11.94 4.32
@@
MedULS + LLaVA-Med 12.21 3.30 10.08 2.42 0.17 27.81 78.85 10.72 11.57 9.06 1.34
LiSA + LLaVA-Med 12.45 0.10 2.18 0.48 0.24 5.05 28.04 2.47 2.99 2.67 0.27
BiomedParse + LLaVA-Med 14.41 0.29 2.25 0.45 0.22 5.46 27.92 2.49 3.02 2.65 0.27
OmniAbnorm-CT ✓ 27.74 20.29 12.51 4.10 1.22 36.56 82.98 13.61 15.83 12.36 5.72
@@
MedULS + LLaVA-Med 11.61 3.35 10.07 2.60 0.32 26.87 78.33 10.61 11.83 9.16 1.79
LiSA + LLaVA-Med 11.36 0.36 2.06 0.43 0.29 5.05 24.51 2.23 2.66 2.15 0.34
BiomedParse + LLaVA-Med 14.25 0.27 2.28 0.66 0.31 5.95 29.05 2.71 3.28 2.81 0.36
OmniAbnorm-CT ✓ 29.89 18.37 12.80 4.51 1.42 36.04 83.97 14.02 16.20 12.79 4.85